A farmer has a rectangular field that measures 100 feet by 150 feet.
He plans to increase the area by 20%. He will do this by increasing
each dimension by the same amount 'x'. What is the equation that represents
the equation of the new field?

Answer by stanbon(75887)

A farmer has a rectangular field that measures 100 feet by 150 feet.
He plans to increase the area by 20%. He will do this by increasing
each dimension by the same amount 'x'. What is the equation that represents
the equation of the new field?
----------------
Ans: (100+x)(150+x) = 1.2*100*150
-----
x^2 + 250x + 15000 = 1.2*15000
------
x^2 + 250x - 0.2*15000 = 0
============================
Cheers.
Stan H.
-------------